### v1

@@ 4.2 Mean Number of Distinct Visited Sites
An alternate measure of exploration efficiency is the MNDVS, which has attracted lots of attention in the diffusive limit and has been studied on lattices with different coordination numbers montroll1969random ; zumofen1982energy . Somewhat surprisingly, this quantity has been studied to a much lesser extent when the movement is persistent with some results for the square lattice appearing only recently larralde2020first .
@@
It is well-known that the MNDVS is related to the first-passage probability via a summation of the first-passage over all states up to some time $t$ rws_on_latticesII ; biroli2022number , which in our case is written as
@@
Via the celebrated renewal equation, the first-passage probability is known to be related to a simpler quantity, the occupation probability rws_on_latticesII ; redner2001guide . In the case of the correlated random walk, if the condition $\widetilde{P}_{\bm{n},\mathfrak{n}}^{(\mathfrak{n})}(\bm{n},z)=\widetilde{P}_{\bm{0},\mathfrak{p}}^{(\mathfrak{p})}(\bm{0},z)$ is met then the renewal equation reduces to the Markovian case larralde2020first ; marris2024persistent . Here this condition is satisfied allowing us to express the first-passage to a postive site as $\widetilde{F}_{\bm{n_{0}},m_{0}}^{(\mathfrak{p})}(\bm{n},z)=\widetilde{P}_{\bm{n_{0}},m_{0}}^{(\mathfrak{p})}(\bm{n},z)/\widetilde{P}_{\bm{n},\mathfrak{p}}^{(\mathfrak{p})}(\bm{n},z)$. Passing this relation into the generating function of Eq. (11) and again making the harmless assumption that the walker started in a positive state, we find
@@
Equation (12) represents the general formulation of the MNDVS for multi-state Markov walks, valid both for unbounded or bounded domains. If the aforementioned condition is met ($\widetilde{P}_{\bm{n},\mathfrak{n}}^{(\mathfrak{n})}(\bm{n},z)=\widetilde{P}_{\bm{0},\mathfrak{p}}^{(\mathfrak{p})}(\bm{0},z)$), i.e. the probability of a walker occupying its initial site is invariant to the location and site type of the initial condition, we may simplify Eq. (12) further to
@@
where we have used the property that the sum of the occupation probability over all the sites and states equals unity to recover the well-known form of the MNDVS in the second line rws_on_latticesII ; blumen1981energy . For non isotropic lattices, e.g. in the presence of reflecting domains, one needs to employ Eq. (12) contrary to what erroneously implied in ref. LucaPRX .
Upon substitution of the generating function of Eq. (5) into Eq. (13) and perfoming some algebra, one arrives at
$\mathcal{M}(t)=\frac{2\pi}{i}\!\oint_{|z|<1}\!\frac{1}{(1\!-\!z)^{2}}\left[\sum_{j=1}^{3}\left(\int_{-\pi}^{\pi}\int_{-\pi}^{\pi}\left[\mathbb{I}-z\bm{\lambda}(\xi_{1},\xi_{2})\right]^{-1}\cdot\bm{U}_{m_{0}}\text{d}\xi_{1}\text{d}\xi_{2}\right)_{j}\right]^{-1}\text{d}z.$ (14)
@@
Figure 3: A comparison between the scout and recruit spreading statistics over the infinite honeycomb lattice with the mean squared displacement in (a) and the mean number of visited sites in (b). The estimated model parameters from data (i.e. $r$, $l$, $b$ and $c$) for both scouts and recruiters are given explicitly in Sec. 2. We display the curves with these parameters with the label ‘Fit’ in the legend and compare them against the diffusive and directed limits. We take the sojourn probability $c$ to be $c\approx 0.7594$ (scouts) and $c\approx 0.7821$ (recruits), leading to $b=0$, and $r=l=(1-c)/2$ and $r=b=l=(1-c)/3$ for the directed and diffusive cases, respectively. In the inset of panel (b) we plot three limiting cases (with dimensionless time), each with $c=0$. The black dots show excellent agreement with the results obtained iteratively in ref. zumofen1982energy for the diffusive walk, while the zero gradient past $t=6$ in the circular case is to be expected since the circular walker will repeat its journey around one hexagon for all time.
In Fig. 3 we plot the MSD and MNDVS for recruits and scouts, with comparison to diffusive and directed spread. It is seen that the scouts explore the space more efficiently than the recruits and both behavioural types explore quicker than a diffusive walker. These results align with expectations, as animals with frontal and bilateral sensory systems are known to exhibit more persistent directional movement compared to pure random walkers turchin1998quantitative ; Martin2013Behaviour . Similarly, scouts are expected to be more effective explorers than recruits, as the latter play a more critical role during food exploitation after scouts make initial patch discoveries.
@@ 5 Search of $\mathcal{N}$ distinct walkers
We first present general theoretical findings in Sec. 5.1 and 5.1 pertaining to search with $\mathcal{N}$ independent random walkers, while in Sec. 5.2 we consider model predictions pertaining to the ant experiment. Specifically, we consider the fFPT, which is the probability that the first walker in the population reaches any target of a set $\mathcal{T}$ for the first time at time $t$, in a heterogeneous population with individuals displaying different movement characteristics. The exact theoretical development for the fFPT that we present below is particularly relevant now given the recent surge in closed-form expressions for the first-passage and splitting probability generating function in a range of Markovian LucaPRX ; giuggioli2023multi and non-Markovian marris2024persistent walks. From the biological perspective, while in general ants share information in the food gathering process AntEcology_chapter_foraging , the application of the fFPT formalism is germane to our experiments as recent work on A. senilis has elucidated that during the exploration phase, i.e. before any food is found, individual ants move largely independent of each other cristin2024spatiotemporal , justifying the assumption of independent walkers.
@@ 5.1 Theory of first first-passage processes in a heterogeneous population with multiple targets
The probability that none of $\mathcal{N}$ independent random walkers have reached any site in the set of targets $\mathcal{T}=\{\bm{s}_{1},...,\bm{s}_{\mathdutchcal{N}}\}$ is expressed via the survival probability of the entire population $\mathcal{S}_{\bm{n}_{0}}(t;\mathcal{N})=\prod_{k=1}^{\mathcal{N}}\mathcal{S}^{(k)}_{\bm{n}_{0}}(t)$ (see e.g. refs. lawley2020universal ; grebenkov2020single and references therein), where $\mathcal{S}^{(k)}_{\bm{n}_{0}}(t)$ is the survival probability of walker $k$ at time $t$. When the movement statistics of each walker are identical and the walkers start from the same initial position, which represents the nest in our case, one has $\mathcal{S}_{\bm{n}_{0}}(t;\mathcal{N})=\mathcal{S}_{\bm{n}_{0}}(t)^{\mathcal{N}}$. In continuous time, the direct relation of the survival probability and the first-passage probability to any of the targets in the set $\mathcal{T}$, namely $F_{\bm{n}_{0}}(\mathcal{T},t)=-\frac{\text{d}}{\text{d}t}\mathcal{S}_{\bm{n}_{0}}(t)$, leads to the well-known expression weiss1983order
@@
Equation (17) represents the time-dependent first-splitting probability for one or more walker among $\mathcal{N}$, starting at $\bm{n}_{0}$, to reach target $\bm{s}_{i}$ before any other in the set of $\mathdutchcal{N}$ targets. As the event (iib) is excluded from all possible outcomes making up $F_{\bm{n}_{0}}(\bm{s}_{i}|\bm{s}_{\{1,…,\mathdutchcal{N}\}-\{i\}},t;\{1,\mathcal{N}\})$, one needs to keep in mind that for $\mathcal{N}\neq 1$, $\sum_{i=1}^{\mathdutchcal{N}}F_{\bm{n}_{0}}(\bm{s}_{i}|\bm{s}_{\{1,…,\mathdutchcal{N}\}-\{i\}},t;\{1,\mathcal{N}\})\neq F_{\bm{n}_{0}}(\mathcal{T},t;\{1,\mathcal{N}\})$.
We now further extend the notion of search for a heterogeneous population to allow for the different movement statistics of the individuals. We do so by partitioning the population into $h\leq\mathcal{N}$ subsets, where each subset is comprised of $\mathcal{N}_{h}$ walkers with the same movement statistics such that $\sum_{h}\mathcal{N}_{h}=\mathcal{N}$ and $\mathcal{N}_{1}\geq\mathcal{N}_{2}\geq...\geq\mathcal{N}_{h}$. By considering all permutations of possible successful search outcomes, the generalisation of Eq. (16), to a heterogeneous population, denoted by the superscript $\mathfrak{h}$, is given by
$F_{\bm{n}_{0}}^{(\mathfrak{h})}(\mathcal{T},t;\{1,\mathcal{N}\})$ $=\sum_{i=1}^{h}\sum_{k_{i}=1}^{\mathcal{N}_{i}}\binom{\mathcal{N}_{i}}{k_{i}}F_{\bm{n}_{0}}^{(i)}(\mathcal{T},t)^{k_{i}}\prod_{j=1}^{h}\mathcal{S}_{\bm{n}_{0}}^{(j)}(t)^{\mathcal{N}_{j}-\delta_{i,j}k_{j}}$
$+\sum_{k_{1}=1}^{\mathcal{N}_{1}}\cdots\sum_{k_{h}=1}^{\mathcal{N}_{h}}\prod_{i=1}^{h}\binom{\mathcal{N}_{i}}{k_{i}}F_{\bm{n}_{0}}^{(i)}(\mathcal{T},t)^{k_{i}}\mathcal{S}_{\bm{n}_{0}}^{(i)}(t)^{\mathcal{N}_{i}-k_{i}},$ (18)
which reduces to Eq. (16) when $\mathcal{N}_{1}=\mathcal{N}$ ($\mathcal{N}_{i}=0$, $\forall i>1$) as we define the nested sum in the second term to be zero in this case. This makes sense physically as the first term represents the case of one or more walkers from the same subset reaching the target at the same time (summed over each subset), while the second term represents the cases where a mixed population of walkers reaches the target for the first time. The splitting probability expression with a heterogeneous population may be derived via an analogous procedure to the homogeneous case, i.e. by replacing the first-passage probabilities in Eq. (18) with the splitting probabilities for the type of walkers belonging to the subset population.
@@
Most studies of the fFPT with $\mathcal{N}$ walkers have focused on large homogeneous populations of walkers searching in unbounded space weiss1983order ; lawley2020universal ; lawley2020probabilistic ; meerson2015mortality . Here instead, guided by the observations of A. senilis, we consider a small population of $\mathcal{N}$ ($\mathcal{N}\leq 10$) individuals with unequal movement statistics. As an illustrative example we consider a search process in bounded periodic space and study two cases: a population comprised of both strongly persistent and diffusive walkers on the square lattice, and a population of correlated walkers, representing scouts and recruits, on a honeycomb lattice.
For the square lattice case we take the well-known diffusive periodic propagator (see e.g. refs. montroll1956random ; barry_hughes_book ; LucaPRX ) and the corresponding correlated propagator given by Eq. (14) of ref. marris2024persistent . For correlated motion on the bounded honeycomb lattice we introduce a periodic boundary on Eq. (5) via standard procedures (see e.g., barry_hughes_book ) to find
$\widetilde{\bm{P}}_{\bm{n}_{0},m_{0}}^{(p)}(\bm{n},z)=\frac{1}{N^{2}}\sum_{k_{1}=0}^{N-1}\sum_{k_{2}=0}^{N-1}\text{e}^{-\frac{2\pi i}{N}\bm{k}\cdot[\bm{n}-\bm{n}_{0}]^{\intercal}}\left[\mathbb{I}-z\bm{\lambda}\left(\frac{2\pi k_{1}}{N},\frac{2\pi k_{2}}{N}\right)\right]^{-1}\cdot U_{m_{0}},$ (19)
@@
where $\mathbb{H}(z)_{l,k}=\alpha_{m_{\bm{s}_{k}}}\widetilde{P}_{\bm{s}_{k},m_{\bm{s}_{k}}}(\bm{s}_{l},m_{\bm{s}_{l}},z)$, $\mathbb{H}(z)_{k,k}=\alpha_{m_{\bm{s}_{k}}}\widetilde{P}_{\bm{s}_{k},m_{\bm{s}_{k}}}(\bm{s}_{k},m_{\bm{s}_{k}},z)$ and $\mathbb{H}^{(i)}(\bm{n}_{0},m_{0},z)$ is the same, but with the $i^{\text{th}}$ column replaced with $\mathdutchcal{F}(n_{0},m_{0_{j}},z)=\alpha_{m_{0_{j}}}\left[\widetilde{P}_{\bm{n}_{0},m_{0_{j}}}(\bm{s}_{1},m_{\bm{s}_{1}},z),...,\widetilde{P}_{\bm{n}_{0},m_{0_{j}}}(\bm{s}{N},m_{\bm{s}_{\mathdutchcal{N}}},z)\right]^{\intercal}$. The splitting probability for the Markovian diffusive case may be found in Eq. (11) of ref. giuggioli2022spatio . The notation $\alpha_{m_{0_{j}}}P_{\bm{n}_{0},m_{0_{j}}}(\bm{n},m,t)=\alpha_{m_{0_{j}}}\bm{e}_{m}^{\intercal}\cdot\bm{P}_{\bm{n}_{0}}(\bm{n},t)\cdot\bm{e}_{m_{0_{j}}}$, is the occupation probability that the walker occupies site $\bm{n}$ and state $m$ given the initial condition $(\bm{n}_{0},m_{0_{j}})$, which is weighted by $\alpha_{m_{0_{j}}}$. Presently, due to the assumption of each direction being equi-probable at time zero $\alpha_{i}=1/3$, $\forall i$. Since each site can be accessed via three internal states, to represent a fully absorbing trap we set the probability of leaving each internal state becomes zero. For example, to place a target in a positive site, absorbing traps in its states $1,2$ and $3$ are required. Hence, for all cases presented here, and elsewhere in this chapter, the entire target set $\mathcal{T}$ is comprised of the number of internal states within the target site. Henceforth, for ease of notation, when discussing specific cases we denote the set of targets with only the physical location(s) at site(s) $\bm{s}_{i}$.
We plot in Fig. 4 the fFPT for different heterogeneous populations on the square lattice. To present results on the square lattice it is convenient to introduce the so-called persistence length tejedor2012optimizing , that is the average number of steps taken in the same direction between turns. On the square lattice it is defined as $p_{l}=[b+2\tau]^{-1}$, where $b$ is the probability of backtracking and $2\tau$ the probability of turning left or right (probability $\tau$ in each direction). In Fig. 4(a) we consider a heterogeneous population of eight walkers searching on a domain of $20\times 20$ lattice sites. The domain contains two targets, one close-by to the initial position and the other far away and just off the persistence path, (see caption of Fig. 4 for the exact locations). By increasing the proportion of persistent walkers in the population (inset Fig. 4(a)) the overall probability $\mathbb{P}$ of finding the far target increases until it is almost equi-probable as many trajectories head in a staight-line out of the initial condition bypassing the closer target.
@@
To gain further insights on the role of persistence in the first-passage processes we interpret our findings also in terms of the probability averages, for which it is known that there is an optimal persistence length $p_{l}$, whose value depends on the search parameters, that minimises the global search time on the square lattice tejedor2012optimizing . If the persistence parameter greatly surpass $p_{l}$, one may find very high mean first-passage time if the target is off the persistence path. Here in Fig. 4(b) we show persistence lengths starting from just past the optimum for the chosen system ($p_{l}\approx 4.76$). In such a parameter regime, an increase in the persistence length hinders the search and weakens the effect of adding more persistent walkers until a population of diffusive walkers outperforms a population of correlated walkers. Around the level of persistence which tips the walker beyond the diffusive efficacy (shown presently as $p_{l}=28.57$), the mean fFPT (MfFPT) becomes non-monotonic as a function of the number of correlated walkers in the population as shown in the inset of panel (b).
For the experimentally relevant case we consider scouts and recruits in a honeycomb lattice and we plot the first-passage dynamics in Fig. 5. In panel (a) we show the splitting probability to a nearby and far target. In such a case, the closest target is always more likely to be found regardless of the proportion of scouts to recruits, which is not only expected to be more energetically efficient, but it also matches the exploratory patterns observed in different ant species martin2024 ; pol2024liquid . Panel (b) shows the entire fFPT distribution for each ratio of scouts to recruits. Since an increase in the proportion of scouts allows the space to be explored more efficiently the so-called direct trajectories become more effective causing the mode of the fFPT distribution to shift upwards and towards the left as it becomes more probable to find the target at shorter times.

### v2

@@ 4.2 Mean Number of Distinct Visited Sites
An alternate measure of exploration efficiency is the MNDVS, which has attracted lots of attention in the diffusive limit and has been studied on lattices with different numbers of neighbouring sites, the so-called coordination number montroll1969random ; zumofen1982energy . Somewhat surprisingly, this quantity has been studied to a much lesser extent when the movement is persistent with some results for the square lattice appearing only recently larralde2020first .
@@
It is well-known that the MNDVS is related to the first-passage probability via a summation of the first-passage over all states up to some time $t$ biroli2022number ; rws_on_latticesII , which in our case is written as
@@
Via the celebrated renewal equation, the first-passage probability is known to be related to a simpler quantity, the occupation probability rws_on_latticesII ; redner2001guide . In the case of the correlated random walk, if the condition $\widetilde{P}_{\bm{n},\mathfrak{n}}^{(\mathfrak{n})}(\bm{n},z)=\widetilde{P}_{\bm{0},\mathfrak{p}}^{(\mathfrak{p})}(\bm{0},z)$ is met then the renewal equation reduces to the Markovian case larralde2020first ; marris2024persistent . Here this condition is satisfied allowing us to express the first-passage to a positive site as $\widetilde{F}_{\bm{n_{0}},m_{0}}^{(\mathfrak{p})}(\bm{n},z)=\widetilde{P}_{\bm{n_{0}},m_{0}}^{(\mathfrak{p})}(\bm{n},z)/\widetilde{P}_{\bm{n},\mathfrak{p}}^{(\mathfrak{p})}(\bm{n},z)$. Passing this relation into the generating function of Eq. (11) and again making the harmless assumption that the walker started in a positive state, we find
@@
Equation (12) represents the general formulation of the MNDVS for multi-state Markov walks, valid both for unbounded or bounded domains. If the aforementioned condition is met ($\widetilde{P}_{\bm{n},\mathfrak{n}}^{(\mathfrak{n})}(\bm{n},z)=\widetilde{P}_{\bm{0},\mathfrak{p}}^{(\mathfrak{p})}(\bm{0},z)$), i.e., the probability of a walker occupying its initial site is invariant to the location and site type of the initial condition, we may simplify Eq. (12) further to
@@
where we have used the property that the sum of the occupation probability over all the sites and states equals unity to recover the well-known form of the MNDVS in the second line blumen1981energy ; rws_on_latticesII . For non-isotropic lattices, e.g., in the presence of reflecting domains, one needs to employ Eq. (12) contrary to what is erroneously implied in ref. LucaPRX .
Upon substitution of the generating function of Eq. (5) into Eq. (13) and performing some algebra, one arrives at
$\mathcal{M}(t)=\frac{2\pi}{i}\!\oint_{|z|<1}\!\frac{1}{(1\!-\!z)^{2}}\left[\sum_{j=1}^{3}\left(\int_{-\pi}^{\pi}\int_{-\pi}^{\pi}\left[\mathbb{I}-z\bm{\lambda}(\xi_{1},\xi_{2})\right]^{-1}\cdot\bm{U}_{\mathfrak{p}}\text{d}\xi_{1}\text{d}\xi_{2}\right)_{j}\right]^{-1}\text{d}z.$ (14)
@@
Figure 3: A comparison between the scout and recruit spreading statistics over the infinite honeycomb lattice with the mean squared displacement in (a) and the mean number of visited sites in (b). The estimated model parameters from data (i.e., $r$, $l$, $b$ and $c$) for both scouts and recruiters are given explicitly in Sec. 2. We display the curves with these parameters with the label ‘Fit’ in the legend and compare them against the diffusive and directed limits. We take the sojourn probability $c$ to be $c\approx 0.7594$ (scouts) and $c\approx 0.7821$ (recruits), leading to $b=0$, and $r=l=(1-c)/2$ and $r=b=l=(1-c)/3$ for the directed and diffusive cases, respectively. In the inset of panel (b) we plot three limiting cases (with dimensionless time), each with $c=0$. The black dots show excellent agreement with the results obtained iteratively in ref. zumofen1982energy for the diffusive walk, while the zero gradient past $t=6$ in the circular case is to be expected since the circular walker will repeat its journey around one hexagon for all time.
In Fig. 3 we plot the MSD and MNDVS for recruits and scouts, with comparison to diffusive and directed spread. It is seen that the scouts explore the space more efficiently than the recruits and both behavioural types explore quicker than a diffusive walker. These results align with expectations, as animals with frontal and bilateral sensory systems are known to exhibit more persistent directional movement compared to pure random walkers Martin2013Behaviour ; turchin1998quantitative . Similarly, scouts are expected to be more effective explorers than recruits, as the latter play a more critical role during food exploitation after scouts make initial patch discoveries.
@@ 5 Search of $\mathcal{N}$ distinct walkers
We first present general theoretical findings in Sec. 5.1 and 5.1 pertaining to search with $\mathcal{N}$ independent random walkers, while in Sec. 5.2 we consider model predictions pertaining to the ant experiment. Specifically, we consider the fFPT, which is the probability that the first walker in the population reaches any target of a set $\mathcal{T}$ for the first time at time $t$, in a heterogeneous population with individuals displaying different movement characteristics. The exact theoretical development for the fFPT that we present below is particularly relevant now given the recent surge in closed-form expressions for the first-passage and splitting probability generating function in a range of Markovian LucaPRX ; giuggioli2022spatio ; sarvaharmangiuggioli2023 and non-Markovian marris2024persistent walks. From the biological perspective, while in general ants share information in the food gathering process AntEcology_chapter_foraging , the application of the fFPT formalism is pertinent to our experiments as recent work on A. senilis has elucidated that during the exploration phase, i.e., before any food is found, individual ants move largely independent of each other cristin2024spatiotemporal , justifying the assumption of independent walkers.
@@ 5.1 Theory of first first-passage processes in a heterogeneous population with multiple targets
The probability that none of $\mathcal{N}$ independent random walkers have reached any site in the set of targets $\mathcal{T}=\{\bm{s}_{1},...,\bm{s}_{\mathdutchcal{N}}\}$ is expressed via the survival probability of the entire population $\mathcal{S}_{\bm{n}_{0}}(t;\mathcal{N})=\prod_{j=1}^{\mathcal{N}}\mathcal{S}^{(j)}_{\bm{n}_{0_{j}}}(t)$ (see e.g., refs. grebenkov2020single ; lawley2020universal and references therein), where $\mathcal{S}^{(j)}_{\bm{n}_{0_{j}}}(t)$ is the survival probability of walker $j$ at time $t$ given it was located at $\bm{n}_{0_{j}}$ at $t=0$. When the movement statistics of each walker are identical and the walkers start from the same initial position, which represents the nest in our case, one has $\mathcal{S}_{\bm{n}_{0}}(t;\mathcal{N})=\mathcal{S}_{\bm{n}_{0}}(t)^{\mathcal{N}}$. In continuous time, the direct relation of the survival probability and the first-passage probability to any of the targets in the set $\mathcal{T}$, namely $F_{\bm{n}_{0}}(\mathcal{T},t)=-\frac{\text{d}}{\text{d}t}\mathcal{S}_{\bm{n}_{0}}(t)$, leads to the well-known expression weiss1983order
@@
Equation (17) represents the time-dependent first-splitting probability for one or more walkers among $\mathcal{N}$, starting at $\bm{n}_{0}$, to reach target $\bm{s}_{i}$ before any other in the set of $\mathdutchcal{N}$ targets. As the event (iib) is excluded from all possible outcomes making up $F_{\bm{n}_{0}}(\bm{s}_{i}|\bm{s}_{\{1,…,\mathdutchcal{N}\}-\{i\}},t;\{1,\mathcal{N}\})$, one needs to keep in mind that for $\mathcal{N}\neq 1$, $\sum_{i=1}^{\mathdutchcal{N}}F_{\bm{n}_{0}}(\bm{s}_{i}|\bm{s}_{\{1,…,\mathdutchcal{N}\}-\{i\}},t;\{1,\mathcal{N}\})\neq F_{\bm{n}_{0}}(\mathcal{T},t;\{1,\mathcal{N}\})$.
We now further extend the notion of search for a heterogeneous population to allow for the different movement patterns and/or different initial conditions of the individuals. We do so by partitioning the population into $h\leq\mathcal{N}$ subsets, where each subset is comprised of $\mathcal{N}_{h}$ walkers with the same movement statistics such that $\sum_{h}\mathcal{N}_{h}=\mathcal{N}$ and $\mathcal{N}_{1}\geq\mathcal{N}_{2}\geq...\geq\mathcal{N}_{h}$. By considering all permutations of possible successful search outcomes, the generalisation of Eq. (16), to a heterogeneous population, denoted by the superscript $\mathfrak{h}$, is given by
$F_{\bm{n}_{0}}^{(\mathfrak{h})}(\mathcal{T},t;\{1,\mathcal{N}\})$ $=\sum_{i=1}^{h}\sum_{k_{i}=1}^{\mathcal{N}_{i}}\binom{\mathcal{N}_{i}}{k_{i}}F_{\bm{n}_{0_{i}}}^{(i)}(\mathcal{T},t)^{k_{i}}\prod_{j=1}^{h}\mathcal{S}_{\bm{n}_{0_{j}}}^{(j)}(t)^{\mathcal{N}_{j}-\delta_{i,j}k_{j}}$
$+\sum_{k_{1}=1}^{\mathcal{N}_{1}}\cdots\sum_{k_{h}=1}^{\mathcal{N}_{h}}\prod_{i=1}^{h}\binom{\mathcal{N}_{i}}{k_{i}}F_{\bm{n}_{0_{i}}}^{(i)}(\mathcal{T},t)^{k_{i}}\mathcal{S}_{\bm{n}_{0_{i}}}^{(i)}(t)^{\mathcal{N}_{i}-k_{i}},$ (18)
which reduces to Eq. (16) when $\mathcal{N}_{1}=\mathcal{N}$ ($\mathcal{N}_{i}=0$, $\forall i>1$) as we define the nested sum in the second term to be zero in this case. This makes sense physically as the first term represents the case of one or more walkers from the same subset reaching the target at the same time (summed over each subset), while the second term represents the cases where a mixed population of walkers reaches the target for the first time. The splitting probability expression with a heterogeneous population may be derived via an analogous procedure to the homogeneous case, i.e., by replacing the first-passage probabilities in Eq. (18) with the splitting probabilities for the type of walkers belonging to the subset population.
@@
Most studies of the fFPT with $\mathcal{N}$ walkers have focused on large homogeneous populations of walkers searching in unbounded space lawley2020universal ; lawley2020probabilistic ; meerson2015mortality ; weiss1983order . Here instead, guided by the observations of A. senilis, we consider a small population of $\mathcal{N}$ ($\mathcal{N}\leq 10$) individuals with unequal movement statistics. As an illustrative example we consider a search process in bounded periodic space and study two cases: a population comprised of both strongly persistent and diffusive walkers on the square lattice, and a population of correlated walkers, representing scouts and recruits, on a honeycomb lattice.
For the square lattice case we take the well-known diffusive periodic propagator (see e.g., refs. LucaPRX ; barry_hughes_book ; montroll1956random ) and the corresponding correlated propagator given by Eq. (14) of ref. marris2024persistent . For correlated motion on the bounded honeycomb lattice we introduce a periodic boundary on Eq. (5) via standard procedures (see e.g., barry_hughes_book ) to find
$\widetilde{\bm{P}}_{\bm{n}_{0},m_{0}}^{(p)}(\bm{n},z)=\frac{1}{N^{2}}\sum_{k_{1}=0}^{N-1}\sum_{k_{2}=0}^{N-1}\text{e}^{-\frac{2\pi i}{N}\bm{k}^{\intercal}\cdot[\bm{n}-\bm{n}_{0}]}\left[\mathbb{I}-z\bm{\lambda}\left(\frac{2\pi k_{1}}{N},\frac{2\pi k_{2}}{N}\right)\right]^{-1}\cdot U_{m_{0}},$ (19)
@@
where $\mathbb{H}(z)_{l,k}=\alpha_{m_{\bm{s}_{k}}}\widetilde{P}_{\bm{s}_{k},m_{\bm{s}_{k}}}(\bm{s}_{l},m_{\bm{s}_{l}},z)$, $\mathbb{H}(z)_{k,k}=\alpha_{m_{\bm{s}_{k}}}\widetilde{P}_{\bm{s}_{k},m_{\bm{s}_{k}}}(\bm{s}_{k},m_{\bm{s}_{k}},z)$ and $\mathbb{H}^{(i)}(\bm{n}_{0},m_{0},z)$ is the same, but with the $i^{\text{th}}$ column replaced with $\mathdutchcal{F}(n_{0},m_{0_{j}},z)=\alpha_{m_{0_{j}}}\left[\widetilde{P}_{\bm{n}_{0},m_{0_{j}}}(\bm{s}_{1},m_{\bm{s}_{1}},z),...,\widetilde{P}_{\bm{n}_{0},m_{0_{j}}}(\bm{s}{N},m_{\bm{s}_{\mathdutchcal{N}}},z)\right]^{\intercal}$. The splitting probability for the Markovian diffusive case may be found in Eq. (11) of ref. giuggioli2022spatio . The notation $\alpha_{m_{0_{j}}}P_{\bm{n}_{0},m_{0_{j}}}(\bm{n},m,t)=\alpha_{m_{0_{j}}}\bm{e}_{m}^{\intercal}\cdot\bm{P}_{\bm{n}_{0}}(\bm{n},t)\cdot\bm{e}_{m_{0_{j}}}$, is the occupation probability that the walker occupies site $\bm{n}$ and state $m$ given the initial condition $(\bm{n}_{0},m_{0_{j}})$, which is weighted by $\alpha_{m_{0_{j}}}$. Presently, due to the assumption of each direction being equi-probable at time zero $\alpha_{i}=1/3$, $\forall i$. Moreover, since each site can be accessed via three distinct internal states, to represent a fully absorbing trap we set the probability of leaving each internal state in the trap site to zero. For example, to place a target in a positive site, absorbing traps in the states $1,2$ and $3$ are required. Hence, for all cases presented here, and elsewhere in this chapter, the target set $\mathcal{T}$ is comprised of sets of three target states for each physical site. Henceforth, for ease of notation, when discussing specific cases we denote the set of targets with only the physical location(s) at site(s) $\bm{s}_{i}$.
We plot in Fig. 4 the fFPT for different heterogeneous populations on the square lattice. To present results on the square lattice it is convenient to introduce the so-called persistence length tejedor2012optimizing , that is the average number of steps taken in the same direction between turns. On the square lattice it is defined as $p_{l}=[b+2\tau]^{-1}$, where $b$ is the probability of backtracking and $2\tau$ the probability of turning left or right (probability $\tau$ in each direction). In Fig. 4(a) we consider a heterogeneous population of eight walkers searching on a domain of $20\times 20$ lattice sites. The domain contains two targets, one close-by to the initial position and the other far away and just off the persistence path, (see caption of Fig. 4 for the exact locations). By increasing the proportion of persistent walkers in the population (inset Fig. 4(a)) the overall probability $\mathbb{P}$ of finding the far target increases until it is almost equi-probable as many trajectories head in a straight-line out of the initial condition bypassing the closer target.
@@
To gain further insights on the role of persistence in the first-passage processes we interpret our findings also in terms of the mean fFPT (MfFPT), for which it is known that there is an optimal persistence length $p_{l}$, whose value depends on the search parameters, that minimises the global search time on the square lattice tejedor2012optimizing . If the persistence parameter greatly surpass $p_{l}$, one may find very high mean first-passage time if the target is off the persistence path. Here in Fig. 4(b) we show persistence lengths starting from just past the optimum for the chosen system ($p_{l}\approx 4.76$). In such a parameter regime, an increase in the persistence length hinders the search and weakens the effect of adding more persistent walkers until a population of diffusive walkers outperforms a population of correlated walkers. Around the level of persistence which tips the walker beyond the diffusive efficacy (shown presently as $p_{l}=28.57$), the MfFPT becomes non-monotonic as a function of the number of correlated walkers in the population as shown in the inset of panel (b).
For the experimentally relevant case we consider scouts and recruits in a honeycomb lattice and we plot the first-passage dynamics in Fig. 5. In panel (a) we show the splitting probability to a nearby and far target. In such a case, the closest target is always more likely to be found regardless of the proportion of scouts to recruits, which is not only expected to be more energetically efficient, but it also matches the exploratory patterns observed in different ant species pol2024liquid ; martin2024 . Panel (b) shows the entire fFPT distribution for each ratio of scouts to recruits. Since an increase in the proportion of scouts allows the space to be explored more efficiently the so-called direct trajectories become more effective causing the mode of the fFPT distribution to shift upwards and towards the left as it becomes more probable to find the target at shorter times.